Find mean, median, mean deviation about mean, mean deviation about median for the


following data


Age in Years 20 – 22 22 – 24 24 – 26 26 – 28 28 – 30 30 – 32 32 – 34


No. of Employees 70 90 110 140 130 80 80


15. Find mean deviation from mode and its coefficient for the following data


Class – Interval 14 – 18 18 – 22 22 – 26 26 – 30 30 – 34


Frequencies 5 12 14 12 7

1
Expert's answer
2022-01-02T17:42:38-0500

14.

"n=700"


"mean=\bar{x}=\dfrac{\sum _if_ix_i}{n}=\dfrac{18960}{700}=27.0857""\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 Class & f_i & Mid\ value & f_ix_i & |x_i-\bar{x}| & |x_i-\bar{x}|f_i\\
& & x_i & & & & \\ \hline
 20-22 & 70 & 21 & 1470 & 6.0857 & 426 \\ \hdashline
 22-24 & 90 & 23 & 2070 & 4.0857 & 367.7143 \\ \hdashline
 24-26 & 110 & 25 & 2750 & 2.0857 & 229.4286 \\ \hdashline
 26-28 & 140 & 27 & 3780 & 0.0857 & 12 \\ \hdashline
 28-30 & 130 & 29 & 4030 & 1.9143 & 248.8571 \\ \hdashline
 30-32 & 80 & 31 & 2640 & 3.9143 & 313.1429 \\ \hdashline
 32-34 & 80 & 33 & 2800 & 5.9143 & 473.1429 \\ \hdashline
 & 700 & & 18960 & & 2070.2857 \\ \hdashline
\end{array}"

Mean deviation of Mean


"\delta \bar{x}=\dfrac{\sum _if_i|x_i-\bar{x}|}{n}=\dfrac{2070.2857}{700}=2.9576"

Coefficient of Mean deviation


"\dfrac{\delta \bar{x}}{\bar{x}}=\dfrac{2.9576}{27.0857}=0.1092"




"median=M"

"=\text{value of } (\dfrac{n}{2})^{th}\text{ observation}"

"=\text{value of } (350)^{th}\text{ observation}"

The median class is "26-28."

"L=" lower boundary point of median class "=26"


"n=" Total frequency "=700"


"cf=" Cumulative frequency of the class preceding the median class "=270"


"f=" Frequency of the median class "=140"


"c=" class length of median class "=2"


Median 

"M=L+\dfrac{\dfrac{n}{2}-cf}{f}\cdot c=26+\dfrac{\dfrac{700}{2}-270}{140}\cdot 2"

"=27.1429"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 Class & f_i & Mid\ value & cf & |x_i-M| & |x_i-M|f_i\\
& & x_i & & & & \\ \hline
 20-22 & 70 & 21 & 70 & 6.1429 & 430 \\ \hdashline
 22-24 & 90 & 23 & 160 & 4.1429& 372.8571 \\ \hdashline
 24-26 & 110 & 25 & 270 & 2.1429 & 235.7143 \\ \hdashline
 26-28 & 140 & 27 & 410 & 0.1429 & 20 \\ \hdashline
 28-30 & 130 & 29 & 540 & 1.8571 & 241.4286 \\ \hdashline
 30-32 & 80 & 31 & 620 & 3.8571 & 308.5714 \\ \hdashline
 32-34 & 80 & 33 & 700 & 5.8571& 468.5714 \\ \hdashline
 & 700 & & & & 2077.1429 \\ \hdashline
\end{array}"

Mean deviation of Median


"\delta \bar{x}=\dfrac{\sum _if_i|x_i-M|}{n}=\dfrac{2077.1429}{700}=2.9673"

Coefficient of Mean deviation


"\dfrac{\delta \bar{x}}{\bar{x}}=\dfrac{2.9673}{27.1429}=0.1093"



15.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 Class & f_i & Mid\ value & |x_i-Z| & |x_i-Z|f_i\\
& & x_i & & & \\ \hline
 14-18 & 5 & 16 & 8 & 40 \\ \hdashline
18-22 & 12 & 20 & 4 & 48 \\ \hdashline
 22-26 & 14 & 24 & 0 & 0 \\ \hdashline
 26-30 & 12 & 28 & 4 & 48 \\ \hdashline
 30-34 & 7 & 32 & 8 & 56 \\ \hdashline
 & 50 & & & 192 \\ \hdashline
\end{array}"

To find Mode Class

Here, maximum frequency is "14."


The mode class is "22-26."


"L=" lower boundary point of mode class "=22"


"f_1="  frequency of the mode class "=14"


"f_0="  frequency of the preceding class "=12"


"f_2="  frequency of the succedding class "=12"


"c="  class length of mode class "=4"


"Z=L+(\dfrac{f_1-f_0}{2f_1-f_0-f_2})\cdot c"

"=22+(\dfrac{14-12}{2(14)-12-12})\cdot 4=24"

Mean deviation of Mode


"\delta \bar{x}=\dfrac{\sum _if_i|x_i-Z|}{n}=\dfrac{192}{50}=3.84"

Coefficient of Mean deviation


"\dfrac{\delta \bar{x}}{\bar{x}}=\dfrac{3.84}{24}=0.16"